In February 1955, a paratrooper fell 370 m from an airplane withoutbeing able to open his chute but happened to land in snow,suffering only minor injuries. Assume that his speed at impact was45 m/s (terminal speed), that his mass (including gear) was 65 kg,and that the magnitude of the force on him from the snow was at thesurvivable limit of 1.4 x 105 N. What are (a) the minimum depth ofsnow that would have stopped him safely and (b) the magnitude ofthe impulse on him from the snow?
Answer #1
survivable limit of 1.4 x 105 N
F= 1.4 x 105 N
retardation =a= F/mass of body
               = 1.4 x 105 /65
                =2153.84 m/s2
a) depth of snow , = y
    using newtons 3rd law we get
    v2 -u2 =2 * a *s
where s= y
a= retadation which is evaluate above
u= 0 ,i.e initial velocity
v=45m/s
so we get y,
y= v2 /(2 * a)
=0.47 m in snow
b)
impulse = change in momentum
= m( 45-0)
= 65 * 45 =2925 kg m/s